Swing Arm Interlock verification

PFC to ARGOS swing arm interlocks have been re-established. Test for both sides (DX/SX) have been performed and functionality verified.

DX+SX laser system startup

DX and SX laser system was started up and nominal position have been adjusted to the set points. However, it looks like the nominal positions with closed loop are non perfect and not always stable! To be investigated further.

DX LAS periscope position measurement:

Set point:

Pup x Pup y Field x Field y
L1 555 126 1106 150
L2 338 430 642 826
L3 176 103 280 90
Periscope nominal positions:

x y flexure loop 343->358 x y flexure loop 358->367 x y flexure loop 367->388 x y
1A 173622 13666 169553 9401 no change 1695523 9401

After back to nominal

position. New position taken as new nominal.
168792 9817
1B -134161 -20182 -131314 -15119 -131314 -15119 -130676 -15741
2A 280787 -52588 279396 -56562 279396 -56562 280041 -56482
2B -217811 82565 -218160 85968 -218160 85968 -218647 85818
3A 164050 124 166601 1085 166601 1085 165722 1068
3B -150178 -14807 -152651 -16066 -152651 -16066 -152069 -16000
SX LAS periscope position measurement:

Set point:

Pup x Pup y Field x Field y
L1 536 109 1085 153
L2 322 430 720 893
L3 165 87 270 206

Periscope nominal positions:

x y 0->382 x y 382-399 x y 400->415 x y 0->48 x y
1A 28375 77410 24329 77643 no change 24329 77643

After back to nominal

position. New position taken as new nominal.
24249 77265 restarted controller 24249 77265
1B -25519 -115566 -23599 -117703 -23599 -117703 -23373 -117962 -23373 -117962
2A -52763 40593 -54724 37133 -54724 37133 -55122 37236 -56372 37764
2B 37085 -69491 37206 -66782 37206 -66782 37552 -66896 38463 -67761
3A -57120 116624 -54481 113912 -54481 113912 -54363 114111 -54363 114111
3B 39573 -202762 37355 -199762 37355 -199762 37289 -200094 37289 -200094
Laser power check:

SX DX DX
L1 9.6 15.0 14.9
L2 10.5 17.4 17.4
L3 8.4 14.5 14.5
SX value with dead and old power meter.

LgswUnitSx caibration with AdSec

From 8:30 to 10:00 the CalUnit interlock is re-enabled.

HEXAPOD 3.77 -1.09 8.00 332 -184
We start to calibrate 10 modes on SX: we load a zero.fits as recontructor.

Modal history must be updated to work with KL_v18 (fits header has been modified by hand, "MH.M2CMATRIX"). Output saved on 10modes_calibrated2.fits.

PBL with fiber ports: we must use the bent gregorian BACK.

# modes IM REC IM LGSW REC COMBINED REC
10 - 20151209_185342 20151209_190000 -
10 20151209_190000 20151209_190714 20151209_191000 20151209_192300
We now close with 20151209_192300 to better recenter the 3 LGS ob their FoV.

PUP X Y
Blue 325 421
Yellow 376 438
Red 487 432
HEXAPOD 3.77 -1.09 8.00 331 -178
# modes IM REC IM LGSW REC COMBINED REC
10 20151209_190000 20151209_202934 20151209_203000 -
50 20151209_203000 20151209_203631 20151209_203800 -
50 20151209_203800 20151209_204641 20151209_205000 -
All IM have been shifted by 4, all have been acquired at 970Hz. Probable PBL with modal amplitudes that seem very low. We multiply the modal history by a factor 2.

New modal amplitudes: amp_calibrated.fits.

The TT signal seems mixed with other signals. We try to slow down the acquisition rate to avoid hitting the AdSec setling time. We set 600Hz as loop speed.

# modes IM REC IM LGSW REC COMBINED REC
50 20151209_205000 20151209_205343 20151209_210500 -
150 - 20151209_211401 20151209_211600 -
Then we crash the shell while acquiring the next IM so we go back to the original modal amplitudes.

# modes IM REC IM LGSW REC COMBINED REC
150 20151209_211600 20151209_211825 20151209_212000 20151209_213200
Blue Pockels cell seems to have the double hear pattern, we check PCT using CalUnit as source, all of them show double hears:

TNs Plot
SX: '20151209_212410','20151209_212541' 20151209_212410_pcTransmission.jpg

Jitter initial position: 20151209_215612

We try again to record a 150 modes IM:

# modes IM REC IM LGSW REC COMBINED REC
150 20151209_212000 20151209_220348 20151209_220500 -

We try also to record a new IM using the x2 multiplied modal amplitudes:

# modes IM REC IM LGSW REC COMBINED REC
150 20151209_220500 20151209_220806 - bad! the signals rms does not increase
150 20151209_220500 20151209_221155 - JOL, x1 modal amplitudes
150 20151209_220500 20151209_221605 - 20151209_221700
150 20151209_221700 20151209_221932 20151209_222000 -

We try to close the loop w/o disturb. Red pupil shows a bad behaviour.

TN Modalplot
20151209_ 222734 -

-- WolfgangGaessler - 11 Dec 2015

Topic revision: r2 - 07 Jan 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback